# SelfDefend: LLMs Can Defend Themselves against Jailbreaking in a Practical Manner

### v1

@@ Abstract
Jailbreaking is an emerging adversarial attack that bypasses the safety alignment deployed in off-the-shelf large language models (LLMs) and has evolved into four major categories: optimization-based attacks such as the Greedy Coordinate Gradient (GCG) attack, jailbreak template-based attacks such as “Do-Anything-Now” (DAN), advanced indirect attacks like DrAttack, and multilingual jailbreaks. However, delivering a practical jailbreak defense is challenging because it needs to not only handle all the above jailbreak attacks but also incur negligible delay to user prompts, as well as be compatible with both open-source and closed-source LLMs.
Inspired by how the traditional security concept of shadow stacks defends against memory overflow attacks, this paper introduces a generic LLM jailbreak defense framework called SelfDefend, which establishes a shadow LLM defense instance to concurrently protect the target LLM instance in the normal stack and collaborate with it for checkpoint-based access control. The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful prompts or intentions in user queries, which we empirically validate using the commonly used GPT-3.5/4 models across all major jailbreak attacks. Our measurements show that SelfDefend enables GPT-3.5 to suppress the attack success rate (ASR) by 8.97-95.74% (average: 60%) and GPT-4 by even 36.36-100% (average: 83%), while incurring negligible effects on normal queries. To further improve the defense’s robustness and minimize costs, we employ a data distillation approach to tune dedicated open-source defense models. These models outperform four SOTA defenses and match the performance of GPT-4-based SelfDefend, with significantly lower extra delays. We also empirically show that the tuned models are robust to targeted GCG and prompt injection attacks.
@@ 1 Introduction
Recent years have witnessed the significant potential of large language models (LLMs) in various domains [2], such as natural language processing (NLP) [3, 4, 5], information retrieval [6], image generation [7], science [8, 9, 10, 11], code tasks [12, 13, 14, 15], security tasks [16, 17, 18, 19, 20, 21], and more. To avoid causing social anxiety, ethical, and legal issues due to LLM responses to harmful questions, LLM vendors typically conduct safety alignment to prevent the misuse of LLMs through techniques like RLHF (Reinforcement Learning from Human Feedback) [22]. In response to a harmful prompt that violates safety policies, an aligned LLM often replies with a standard response such as “I’m sorry, I can’t assist with that request.” To bypass LLMs’ safety alignment checks, an adversarial attack known as jailbreaking [23] was proposed.
@@
In the past two years, research on LLM jailbreak attacks and defenses has attracted considerable interest, with most of them focused on the offensive side. Specifically, jailbreak strategies have evolved from manual prompt engineering [23, 24, 25] to automatic LLM-based red teaming [26, 27, 28]. Besides these template-based jailbreaks aimed at identifying a valid jailbreak prompt template, a more generic, optimization-based adversarial jailbreak approach, notably Greedy Coordinate Gradient (GCG) [29], was recently proposed. It uses public available models to train adversarial suffixes that maximize the probability for LLMs to produce an affirmative response rather than refusing to answer. These studies [29, 30, 31] found that the identified suffixes are transferable to closed-source off-the-shelf LLMs. More recently, advanced indirect jailbreaks like DrAttack [32] and Puzzler [33], as well as multilingual jailbreaks [34, 35], have also been invented. In addition to proposing these four major types of jailbreaks, various benchmark studies on LLM jailbreak attacks [23, 36, 37, 38, 39, 40] have also been conducted.
@@
On the contrary, the defensive side is somewhat overshadowed, despite around a dozen defense mechanisms being proposed in the past year. They can be roughly categorized into tuning-based and non-tuning-based mechanisms. Specifically, tuning-based defenses [41, 42, 43, 44, 45, 46] aim to fundamentally improve a model’s safety alignment against jailbreaking, while non-tuning-based defenses [47, 48, 49, 50, 51, 52, 53, 54, 55] can be typically applied to any off-the-shelf LLMs. We conduct an analysis of these defense techniques in §3 and find that it is still challenging for them to be widely used in practice. In short, we advocate that a practical jailbreak defense needs to not only handle all the aforementioned jailbreak attacks but also incur negligible delay to user prompts, as well as be compatible with both open-source and closed-source LLMs in various scenarios (see details in §3).
@@
In this paper, we propose a new perspective on practically defending against jailbreak attacks, inspired by how the traditional security concept of shadow stacks [56] defends against memory overflow attacks. Similar to the shadow stack creating a shadow memory space, we establish a shadow LLM defense instance, $LLM_{defense}$, alongside the target LLM instance, $LLM_{target}$, in the normal stack. Under this framework, $LLM_{target}$ can process any user prompt query $P_{query}$ as usual to produce a token-by-token output. Meanwhile, $LLM_{defense}$ employs a dedicated detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful prompt (via $P_{direct}$) or intention (via $P_{intent}$). Such a unique setup brings several benefits. ① It simultaneously utilizes both $LLM_{target}$’s own safety alignment and $LLM_{defense}$’s dedicated jailbreak detection, largely increasing the defense success rate due to this dual-layer protection. ② Since $LLM_{defense}$’s output is typically short, such as only “No” (indicating no issue) for normal queries, a checkpoint in the normal stack tends to be quickly triggered from the shadow stack without delaying $LLM_{target}$’s output. ③ Since $LLM_{defense}$ does not need to modify or monitor $LLM_{target}$’s internals, it can protect both open-source and closed-source LLMs. We concretize the above ideas into a generic jailbreak defense framework called SelfDefend and will introduce its details in §4.
@@
The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful portions (prompts/intentions) in user queries. To validate this hypothesis, we conduct an empirical measurement in §5 using the commonly used GPT-3.5/4 models under the SelfDefend architecture to test all major jailbreak attacks [29, 31, 36, 28, 32, 33, 34]. The results are quite promising in that SelfDefend enables both GPT-3.5 and GPT-4 to significantly suppress the attack success rate (ASR). Specifically, GPT-3.5-based SelfDefend reduces the ASR by 8.97-95.74% (average: 60%) compared to the baseline GPT-3.5, lowering the ASR to an average of 0.28, and GPT-4-based SelfDefend even reduces the ASR by 36.36-100% (average: 83%) compared to the baseline GPT-4, lowering the ASR to an extremely low average of 0.06. Besides the jailbreak scenarios, we also test SelfDefend against 805 normal prompts from the AlpacaEval dataset [57] and find that the pass rate is not affected for GPT-3.5 and slightly decreases by 3% for GPT-4, indicating that SelfDefend incurs negligible effects on normal queries. Moreover, SelfDefend incurs zero delay for over 95% of normal prompts across three out of four configurations, averaging 0.08 seconds for GPT-3.5 and 0.04 seconds for GPT-4. In various attack scenarios, SelfDefend causes an overall average extra delay of 0.04 seconds for GPT-3.5 and 0.39 seconds for GPT-4, respectively.
@@
While the above measurements indicate that GPT-based SelfDefend effectively reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.06, GPT-4 itself is commonly known to be expensive [58]. Moreover, the closed-source nature of GPT-3.5/4 raises privacy concerns for non-OpenAI model vendors. Therefore, we attempt to tune an open-source model that can be used under the SelfDefend framework for robust, low-cost, and self-contained jailbreak defense. By conducting GPT-4-based data distillation (with SelfDefend’s prompts) on a red-team dataset from Anthropic [59] comprising 38,961 harmful and harmless prompts, we generate a large set of high-quality tuning data. The distilled datasets are then used to tune our defense models through the LoRA fine-tuning technique [60] on the publicly available Llama-2-7b model [61]. The details will be introduced in §6.
@@
To extensively evaluate our tuned models, we not only assess SelfDefend’s performance as in the aforementioned empirical measurement, but also compare with four representative jailbreak defenses: Perplexity Filter [62], SmoothLLM [51], Llama Guard [42], and Llama Guard 2 [63]. The results show that SelfDefend consistently outperforms all other defenses in 19 out of 21 tested attack scenarios (seven jailbreak methods and three target LLMs: GPT-3.5, GPT-4, and Llama 2) and reaches the level of GPT-4-based SelfDefend’s defense. The only two exceptions occurred in comparison with Perplexity Filter [62], which is specifically designed to defend against the GCG attack and offers no defense against other types of jailbreaks, particularly for the GCG attack targeting GPT-3.5 and Llama 2 LLMs. Nevertheless, we still outperform it when the target LLM is GPT-4.
@@
Besides defense effectiveness, we also measure the extra delay $\Delta d$ caused by our defense models and find that the average $\Delta d$ is negligible, at 0-0.01 seconds for normal prompts. For attack scenarios, the maximum $\Delta d$ has decreased from 1.1 seconds in GPT-4-based SelfDefend to 0.25 seconds, with $\Delta d$ in all attack scenarios except for JSAA now below 0.1 seconds. These findings indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it feasible for real-world deployment. Furthermore, we specifically assess whether the detected harmful portion actually aligns with the original prompt through the ensemble CLIP-score [64], and empirically show that the tuned models are robust to targeted GCG and prompt injection [65, 66] attacks that have white-box access to our defense models. More details are available in §7.
Availability and Ethics. To facilitate future research, we are on the way of releasing our datasets, raw evaluation results, and code at this anonymous GitHub link: https://github.com/SelfDefend. While our study required testing various jailbreak attacks, these tests never affected other OpenAI users and were conducted solely on our own OpenAI account and a local server. We also tried to minimize the overhead to OpenAI by spreading all the experiments over a period of two months instead of just several days and paid around 1,500 USD to OpenAI for our API usage.
@@ 2.1 Threat Model
Attack Scenario. This research focuses on the attack scenario where an adversary seeks to perform jailbreaking on a text-based large language model (LLM). Multimodal jailbreaks [67, 68, 69, 70] are outside the scope of this paper. The objective of jailbreaking is to bypass the LLM’s safety mechanisms and induce it to generate harmful, toxic, or objectionable content. In this context, we assume the adversary has access to the LLM and can input arbitrary prompts to it. Given the vocabulary $\mathcal{T}$, the LLM, denoted as $\text{LLM}:\mathcal{T}^{\star}\to\Delta(\mathcal{T})$, takes a sequence of tokens $\mathcal{T}^{\star}$ as input and outputs a distribution $\Delta(\mathcal{T})$ over the next token in the sequence.
@@
Adversary’s Objective. The adversary’s main objective is to generate responses from the LLM that are classified as successful jailbreaks. Specifically, the adversary seeks to maximize the probability that a response $R$ generated by the LLM for a given prompt $P$ is classified as harmful according to the goal $G$. Formally, the adversary’s objective can be expressed as:
@@
where the randomness is due to the stochastic nature of the LLM’s responses to the input prompt $P$. In essence, the adversary iterates over potential prompts to find one that maximizes the likelihood of producing a harmful output as judged by the classifier.
Constraints and Assumptions. We assume the following constraints and assumptions in the general jailbreak threat model:
  • The adversary has black-box access to the LLM, meaning they can input prompts and observe outputs but do not have direct access to the model’s internal parameters or training data.
  • The goal string $G$ is predefined and represents a specific type of harmful content that the adversary aims to induce from the LLM.
  • The classifier JUDGE is accurate in determining whether a response constitutes a successful jailbreak according to the harmful goal $G$.
Under these assumptions, the adversary’s challenge is to discover prompts that effectively exploit the LLM’s vulnerabilities, leading to the generation of undesirable content while navigating the constraints of black-box access.
@@ 2.2 Jailbreak Attacks
Existing jailbreak attacks can be roughly grouped into four categories: optimization-based, template-based, indirect, and multilingual jailbreaks.
@@
Optimization-based Jailbreak typically appends a suffix to a malicious prompt and iteratively updates this suffix using gradient-based or search-based methodologies for jailbreaking. The pioneering work by GCG [29] introduced a method known as the greedy coordinate gradient to optimize adversarial suffixes, facilitating transferable jailbreaks across various prompts and models. Sitawarin et al. [30] further extended this technique to GCG++ by employing a proxy model to guide the optimization process. Beyond gradient-based optimization, Andriushchenko et al. [31] utilized a simple random search on a suffix to maximize the target probability.
Template-based Jailbreak involves the use of prompt engineering [23, 24, 25] or LLM-based red teaming [26, 27, 36, 71, 28] to create effective prompt templates that can mislead LLMs into producing restricted content. Early template-based attacks combine manually crafted templates [23, 24, 25] with malicious instructions to achieve jailbreaking. For example, Liu et al. [24] proposed role-playing prompts to confuse LLMs. A more sophisticated approach, AutoDAN [28], employs genetic algorithms to optimize these hand-crafted templates. Recent advancements include the use of LLMs to assist in generating jailbreak prompts. PAIR [36] is the first method to utilize an auxiliary LLM to generate jailbreaks. Mehrotra et al. [71] then refined PAIR’s approach through a tree-based search method.
@@
Indirect Jailbreak conceals malicious intents within the query text to circumvent the safety mechanisms of LLMs and elicit the desired malicious response. A straightforward method for executing indirect jailbreaks is by performing word substitution on the original malicious instruction [72]. Recently, DrAttack [32] introduced a technique that decomposes a malicious prompt into separate sub-prompts, effectively masking its underlying malicious intent. Puzzler [33] provides clues related to the malicious prompt, thereby guiding the target LLM towards a jailbreak.
@@
Multilingual Jailbreak translates the harmful prompt into a language in which LLMs are less aligned for safety [34, 35]. Deng et al. [34] found that it is easier to jailbreak LLMs in low-resource languages than in high-resource languages and released a multilingual jailbreak prompt dataset, MultiJail. Besides the multilingual strategy, Wei et al. [23] encode the original harmful prompt into Base64 format to reduce the sensitivity of LLMs.
@@
TABLE I: A comparison of existing jailbreak defenses under the four objectives we envisioned in §3.1. Note that the first six rows denote tuning-based defenses, while the last seven rows represent non-tuning-based defenses.
@@
Llama Guard [42] arXiv:2312.06674 Align the classification of both prompt and response
@@
DRO [43], PAT [44] SeT @ ICLR’24 Add a safety prompt and optimize it as a prefix using prompt tuning
SafeDecoding [45] arXiv:2402.08983 Compare and amplify the token probabilities of safety disclaimers
@@
RAIN [48] ICLR’24 Self-evaluate each token output, rewind, and determine the final output
[50, 51, 52] arXiv:2309.14348 Perturb copies of each prompt and aggregate their output responses
IAPrompt [53] arXiv:2401.06561 Conduct a chain-of-thought intention analysis for each input prompt
GradSafe [54] arXiv:2402.13494 Compare the prompt’s gradient similarity with safety-critical gradients
  • = applies; = partially applies; = does not apply.
  • 1
    We only present the venue information for the earliest article in each line of defense techniques.

### v2

@@ Abstract
Jailbreaking is an emerging adversarial attack that bypasses the safety alignment deployed in off-the-shelf large language models (LLMs) and has evolved into multiple categories: human-based, optimization-based, generation-based, and the recent indirect and multilingual jailbreaks. However, delivering a practical jailbreak defense is challenging because it needs to not only handle all the above jailbreak attacks but also incur negligible delays to user prompts, as well as be compatible with both open-source and closed-source LLMs.
Inspired by how the traditional security concept of shadow stacks defends against memory overflow attacks, this paper introduces a generic LLM jailbreak defense framework called SelfDefend, which establishes a shadow LLM as a defense instance to concurrently protect the target LLM instance in the normal stack and collaborate with it for checkpoint-based access control. The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful prompts or intentions in user queries, which we empirically validate using the commonly used GPT-3.5/4 models across all major jailbreak attacks. To further improve the defense’s robustness and minimize costs, we employ a data distillation approach to tune dedicated open-source defense models. These models outperform six state-of-the-art defenses and match the performance of GPT-4-based SelfDefend, with significantly lower extra delays. We also empirically show that the tuned models are robust to adaptive jailbreaks and prompt injections.
@@ 1 Introduction
Recent years have witnessed the significant potential of large language models (LLMs) in various domains [90], such as natural language processing (NLP) [48, 94, 38], information retrieval [97], image generation [52], science [24, 83, 57, 68], code tasks [50, 36, 37, 72], security tasks [21, 65, 78, 64, 60, 41], and more. To avoid causing social anxiety, ethical, and legal issues due to LLM responses to harmful questions, LLM vendors typically conduct safety alignment to prevent the misuse of LLMs through techniques like RLHF (Reinforcement Learning from Human Feedback) [30]. In response to a harmful prompt that violates safety policies, an aligned LLM often replies with a standard response such as “I’m sorry, I can’t assist with that request.” To bypass LLMs’ safety alignment, an adversarial attack known as jailbreaking [70] was proposed.
@@
In the past two years, research on LLM jailbreak attacks and defenses has attracted considerable interest, with most of them focused on the offensive side. Jailbreak strategies have evolved from manual prompt engineering [70, 43, 73, 62] to automatic LLM-based red teaming [54, 16]. Besides these human-crafted and generative jailbreaks aimed at identifying a valid jailbreak prompt, a more generic, optimization-based adversarial jailbreak approach, notably Greedy Coordinate Gradient (GCG) [99], was proposed. It learns adversarial suffixes on public-available models to maximize their probability in producing an affirmative response instead of refusing, which can be transferable to closed-source off-the-shelf LLMs. Recently, advanced indirect jailbreaks like DrAttack [33] and Puzzler [10], as well as multilingual jailbreaks [17, 61], have also been invented. In addition to proposing new attacks, various benchmark studies on LLM jailbreak attacks [46, 32, 14, 86] have also been conducted.
@@
On the contrary, the defensive side is somewhat overshadowed, despite over a dozen defense mechanisms being proposed in the past year. They can be roughly categorized into model-based and plugin-based mechanisms. Specifically, model-based defenses [35, 89, 96, 81, 93, 49, 82, 45, 77, 91, 39] aim to fundamentally improve a model’s robustness against jailbreaking, while plugin-based defenses [4, 26, 55, 29, 9, 59, 27, 73, 80, 88, 25, 79, 23, 98] can be typically plugged into any off-the-shelf LLMs. We conduct an analysis of these defense techniques in §3 and find that it is still challenging for them to be widely used in practice. In short, we advocate that a practical jailbreak defense needs to not only handle all the aforementioned jailbreak attacks but also incur negligible delay to user prompts, as well as be compatible with both open-source and closed-source LLMs.
@@
In this paper, we propose a new perspective on defending jailbreak attacks, inspired by how the traditional security concept of shadow stacks [8] defends against memory overflow attacks. Similar to the shadow stack creating a shadow memory space, we establish a shadow LLM defense instance, $LLM_{defense}$, alongside the target LLM instance, $LLM_{target}$, in the normal stack. Under this framework, $LLM_{target}$ can process any user prompt query $P_{query}$ as usual to produce a token-by-token output. Meanwhile, $LLM_{defense}$ employs a tailored detection prompt, $P_{direct}$ or $P_{intent}$, to wrap $P_{query}$ and detect its harmful prompt (via $P_{direct}$) or intention (via $P_{intent}$). Such a unique setup brings several benefits. ① It simultaneously utilizes both $LLM_{target}$’s safety alignment and $LLM_{defense}$’s jailbreak detection, largely increasing the defense success rate due to this dual-layer protection. ② As $LLM_{defense}$’s output is typically short, such as only “No” (indicating no issue) for normal queries, a checkpoint in the normal stack tends to be quickly triggered from the shadow stack without delaying $LLM_{target}$’s output. ③ Since $LLM_{defense}$ does not need to modify or monitor $LLM_{target}$’s internals, it can protect both open-source and closed-source LLMs. We concretize the above ideas into a generic jailbreak defense framework called SelfDefend and will introduce its details in §4.
@@
The effectiveness of SelfDefend builds upon our observation that existing LLMs (both target and defense LLMs) have the capability to identify harmful portions (prompts/intentions) in user queries. To validate this hypothesis, we conduct an empirical measurement in §5 using the commonly used GPT-3.5/4 models under the SelfDefend architecture to test all major jailbreak attacks [62, 99, 40, 12, 47, 33, 10, 17]. The results are quite promising in that SelfDefend enables both GPT-3.5 and GPT-4 to significantly suppress the attack success rate (ASR). Specifically, GPT-3.5-based SelfDefend reduces the ASR by 8.97-97.26% (average: 65.06%) compared to the baseline GPT-3.5, lowering the ASR to an average of 0.244, and GPT-4-based SelfDefend even reduces the ASR by 69.69-100% (average: 85.53%) compared to the baseline GPT-4, lowering the ASR to an extremely low average of 0.063. Besides the jailbreak scenarios, we also test SelfDefend against 805 normal prompts from the AlpacaEval dataset [34] and find that the pass rate is almost unaffected for GPT-3.5 and slightly decreases by 2.77% for GPT-4, indicating that SelfDefend incurs negligible effects on normal queries. Moreover, SelfDefend incurs zero delay for over 95% of normal prompts across three out of four configurations. For multiple jailbreak samples, SelfDefend causes an average extra delay of 0.05 seconds for GPT-3.5 and 0.40 seconds for GPT-4, respectively.
@@
While the above measurements indicate that GPT-based SelfDefend effectively reduces the success rates of various types of jailbreak attacks, lowering the ASR to an average of 0.063, GPT-4 itself is commonly known to be expensive [3]. Moreover, the closed-source nature of GPT-3.5/4 raises privacy concerns for non-OpenAI model vendors. Therefore, we attempt to tune an open-source model that can be used under the SelfDefend framework for robust, low-cost, and self-contained jailbreak defense. By conducting GPT-4-based data distillation (with SelfDefend’s prompts) on a red-team dataset from Anthropic [19] comprising 38,961 harmful and harmless prompts, we generate a large set of high-quality tuning data, which is then used to tune our defense models through LoRA fine-tuning [22] on the publicly available Llama-2-7b model [67]. The details will be introduced in §6.
@@
To extensively evaluate our tuned models, we not only assess SelfDefend’s performance as in the aforementioned empirical measurement, but also compare with six representative jailbreak defenses: ICD [73], SafeDecoding [82], Perplexity Filter [26], SmoothLLM [59], Llama Guard [25], and Llama Guard 2 [66]. The results show that SelfDefend consistently outperforms all other defenses in 20 out of 24 tested attack scenarios (involving eight jailbreak methods and three target LLMs: GPT-3.5, GPT-4, and Llama-2) and reaches the level of GPT-4-based SelfDefend’s defense.
@@
Besides defense effectiveness, we also measure the extra delay $\Delta d$ caused by our defense models and find that the average $\Delta d$ is negligible, at 0-0.01 seconds for normal prompts. For attack scenarios, the maximum $\Delta d$ has decreased from 1.56 seconds in GPT-4-based SelfDefend to 0.27 seconds, with $\Delta d$ in all attack scenarios except for DAN now below 0.1 seconds. These findings indicate that the tuning-based SelfDefend achieves negligible delays for both normal and jailbreak prompts, making it efficient for potential real-world deployment. Furthermore, we specifically assess whether the detected harmful portion actually aligns with the original prompt through the ensemble CLIP-score [58], and empirically show that the tuned models are robust to adaptive attacks and prompt injections [44, 42]. Details are available in §7.
The contributions of this paper are summarized as follows:
  • We creatively apply the traditional system security concept of shadow stacks to practical LLM jailbreak defense, and our SelfDefend framework utilizes LLMs in both normal and shadow stacks for dual-layer protection.
  • We successfully initialize SelfDefend for GPT-3.5/4 with two carefully designed detection prompts and empirically validate that LLMs can identify harmful portions (prompts/intentions) in user queries using our measures.
  • We further fine-tune dedicated open-source models that can be used under the SelfDefend architecture for robust, low-cost, and self-contained jailbreak defense.
@@ 2.1 Threat Model
Attack Scenario. This research focuses on the attack scenario where an adversary seeks to perform jailbreaking on a text-based large language model (LLM). Multimodal jailbreaks [56, 6, 76, 51] are outside the scope of this paper. The objective of jailbreaking is to circumvent the LLM’s safety alignment and induce it to generate harmful, toxic, or objectionable content. In this context, we assume the adversary can access the target LLM’s interface and input arbitrary prompts to it. Given the vocabulary $\mathcal{T}$, the LLM, denoted as $\text{LLM}:\mathcal{T}^{\star}\to\Delta(\mathcal{T})$, takes a sequence of tokens $\mathcal{T}^{\star}$ as input and outputs a distribution $\Delta(\mathcal{T})$ over the next token in the sequence.
@@
Adversary’s Objective. The adversary’s objective is to generate responses from the LLM that are classified as successful jailbreaks. Specifically, the adversary seeks to maximize the probability that a response $R$ generated by the LLM for a given prompt $P$ is classified as harmful according to the goal $G$. Formally, the adversary’s objective can be expressed as:
@@
where $\Pr$ is the probability, and the randomness is due to the stochastic nature of the LLM’s responses to the input prompt $P$. Essentially, the adversary iterates over potential prompts to find one that maximizes the likelihood of producing a harmful output, as judged by the classifier.
Constraints and Assumptions. We assume the following setup that is commonly used in the jailbreak threat model:
  • The adversary requires only black-box access to the LLM, meaning they can input prompts and observe outputs but do not necessarily need access to the model’s internal parameters or training data.
  • The goal string $G$ is predefined and represents a specific type of harmful content that the adversary aims to induce.
  • The classifier JUDGE accurately determines if a response constitutes a successful jailbreak based on $G$.
@@ 2.2 Jailbreak Attacks
Existing jailbreak attacks can be roughly grouped into multiple categories: human-based, optimization-based, generation-based, and the recent indirect and multilingual jailbreaks.
Human-based Jailbreak involves manually crafting jailbreak prompts to exploit LLM vulnerabilities [43, 70, 73, 62, 16]. Wei et al. [70] utilize two jailbroken modes of LLMs (i.e., out-of-distribution inputs and the conflict between the model’s capabilities and safety goals) to guide the design of manual jailbreaks. Deng et al. [16] engineered a proof-of-concept (PoC) jailbreak prompt that alters an LLM’s output to generate harmful content by making it act as AIM (Always Intelligent and Machiavellian) and used it as a seed to create more jailbreak prompts. Additionally, Shen et al. [62] present the JailbreakHub framework, a platform for crowdsourcing jailbreak prompts from online contributors.
@@
Optimization-based Jailbreak typically updates the adversarial prompt iteratively using gradient-based or search-based methodologies [99, 63, 5, 40, 28]. The pioneering work by GCG [99] introduced a method known as the greedy coordinate gradient to optimize adversarial suffixes, facilitating transferable jailbreaks across various prompts and models. Sitawarin et al. [63] further extended this technique to GCG++ by employing a proxy model to direct the optimization process. Beyond gradient-based optimization, Andriushchenko et al. [5] utilized a simple random search on a suffix to increase the likelihood of hitting the target probability. Unlike optimizing these obviously unreadable suffixes, AutoDAN [40] automatically constructs human-readable jailbreak prompts using a carefully designed hierarchical genetic algorithm.
Generation-based Jailbreak employs language models [54, 16, 12, 47, 53] to generate effective jailbreak prompts that can mislead LLMs into producing restricted content. An intuitive approach is to use an auxiliary LLM to construct candidate prompts through prompt engineering. For example, PAIR [12] fed the response of the target model back to the attacking LLM to adapt the output for deceptive jailbreaks. Mehrotra et al. [47] then refined PAIR’s approach through tree-of-thought reasoning [84]. Moreover, the adversary can train a new LLM specifically to attack the target model. For example, Paulus et al. [53] fine-tuned the Advprompter LLM to generate human-readable suffixes against the target LLM.
@@
Indirect Jailbreak conceals malicious intents within the query text to circumvent the safety mechanisms of LLMs and elicit the desired malicious response [20, 33, 10, 70]. A straightforward method for executing indirect jailbreaks is to perform word substitution on the original malicious instruction [20]. Recently, DrAttack [33] introduced a technique that decomposes a malicious prompt into separate sub-prompts, effectively masking its underlying malicious intent. Meanwhile, Puzzler [10] provides clues related to the malicious prompt, thereby inducing the target LLM toward a jailbreak.
@@
Multilingual Jailbreak translates the harmful prompt into a language in which LLMs are less aligned for safety [17, 85, 61, 31, 70, 87]. Deng et al. [17] found that it is easier to jailbreak LLMs in low-resource languages, such as Zulu [85], than in high-resource languages and released a multilingual jailbreak prompt dataset, MultiJail. Besides the multilingual strategy, Wei et al. [70] and Yuan et al. [87] adopted an obfuscation strategy [15] to either encode or encrypt the original harmful prompt, thereby reducing the sensitivity of LLMs.
@@
Table 1: A comparison of existing jailbreak defenses under the four objectives we envisioned in §3.1. Note that the first 11 rows denote plugin-based defenses, while the last nine rows represent model-based defenses.
@@
Smooth [9, 59, 27] ACL 2024 (2309.14348) Perturb copies of each prompt and aggregate their output responses
ICD [73] arXiv:2310.06387 Use safe in-context demonstrations to enhance the model’s robustness
Self-Reminder [80] NMI 2023 (December 2023) Add a self-reminder system prompt to make ChatGPT respond safely
Llama Guard [25] arXiv:2312.06674 An input-output safeguard for safety classification of prompts and response
IAPrompt [88] arXiv:2401.06561 Conduct an intention analysis for each input prompt
GradSafe [79] ACL 2024 (2402.13494) Compare the prompt’s gradient similarity with safety-critical gradients
Circuit Breaking [98] arXiv:2406.04313 Interrupt the model output harmful content at the internal representations
RAIN [35] ICLR 2024 (2309.07124) Self-evaluate each token output, rewind, and determine the final output
Goal Prioritization [89] ACL 2024 (2311.09096) Integrate instructions to control the priority between helpfulness and safety
@@
DRO [93], PAT [49] ICML 2024 (2401.18018) Add a prefix as safety prompt and optimize it with prompt tuning
SafeDecoding [82] ACL 2024 (2402.08983) Compare and amplify the token probabilities of safety disclaimers
Eraser [45] arXiv:2404.05880 Encourage LLMs to forget harmful knowledge via machine unlearning
CAT [77] arXiv:2405.15589 Conduct adversarial training on continuous embedding attacks
@@
Adversarial Tuning [39] arXiv:2406.06622 Fine-tune the LLM with token- and semantic-level adversarial prompts
  • = applies; = partially applies; = does not apply. Note that we only present the venue information for the earliest article in each line of defense techniques.